GlueX TOF Meeting, July 3, 2014

From GlueXWiki
Jump to: navigation, search

GlueX Time-of-Flight Meeting
Thursday, July 3, 2014
10:00 am EDT
JLab: CEBAF Center, F326


  1. Announcements
    1. New BlueJeans Meeting ID for TOF: 350531998
  2. TOF Trigger (Sascha S.)
  3. Review of June 19th Meeting
  4. Installation Update (Hussein)


Enter BlueJeans Meeting
  1. To join via Polycom room system go to the IP Address: ( and enter the meeting ID: 350531998
  2. To join via a Web Browser, go to the page
  3. To join via phone, use one of the following numbers and the meeting ID: 350531998
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560


Talks can be deposited in the directory /group/halld/www/halldweb/html/talks/2014-3Q on the JLab CUE. This directory is accessible from the web at .



  • FSU: Sasha Ostrovidov, Aristeidis Tsaris
  • JLab: Hussein Al Ghoul, Mark Ito (chair), Sasha Somov

Trigger and Data Acquisition

Sascha S. and Hussein gave us a report.

  • The trigger output from each board was checked: flash-to-CTP route is working.
  • Thresholds [?] were checked for all channels.
  • All PMTs were set at 1250 V, on the conservative side.
  • A pedestal calibration was done on the flashes and the DACs were set to give a pedestals of 200 counts.
  • Pedestals widths are good, about 1.3 counts rms.
  • Two trigger methods were tested:
    1. instantaneous energy (single-sample ADC count over threshold)
    2. integrated energy over an 80 ns window (20 samples)
  • There were some initial problems with the CAEN TDCs that had to be worked through.
  • Data was written to disk with both flash and TDC data with the TOF trigger.
  • A stand-alone analyzer was written to create ROOT trees directly from the EVIO data.
  • Sascha showed an event (run tof_106.0, event 1)

The first figure is a graphical representation showing one counter hit in the vertical plane and two hit in the horizontal. The next four figures show the flash data from each channel of each of the four ends of the detectors (up, down, north, and south). In particular the two horizontal counters show the expected coincidence in time and similarity in amplitude.

  • The trigger rate was about 200 kHz.
  • There is one bad splitter and one bad cable in the system at present. Chris Stanislav will make the replacement.

Sascha S. will turn will talk to Hussein later today to start turning over the data acquisition to the FSU team. Aristeidis developed an independent viewer and had started looking at the data. Paul Eugenio is working on getting scaler information going, as a monitor on light leaks among other things.

Now that the DAQ is working at some level there is a huge list of commissioning issues that can be addressed. We decided to try and have a meeting on Monday to make a list and assign priorities. Mark will try to set this up.

BlueJeans Meeting Number

Mark has booked a new meeting in BlueJeans for these bi-weekly meetings. This is only to give him a bit more direct control of meeting details. The new meeting ID is 350531998. The BlueJeans meeting page has been updated.

Additional Plots